### v1

@@ 5.4 Hyperparameter Analysis(RQ3)
Refer to caption (a) ETTh1
Refer to caption (b) ETTh2
Refer to caption (c) ETTm1
Refer to caption (d) Electricity
@@
Refer to caption (a) ETTh1
Refer to caption (b) ETTh2
Refer to caption (c) ETTm1
Refer to caption (d) Electricity
@@ 5.4.1 The Impact of $\lambda$
As our model jointly optimizes the supervised and self-supervised tasks with hyperparameter $\lambda$ in Eq.(14), we first explore the effect of $\lambda$ on the model performance. The value of $\lambda$ determines the importance of the self-supervised task relative to the supervised task, we tune it in $\{0,0.1,0.25,0.33,0.5,0.66,\\ 0.8,0.9,1.0\}$ to find the best balance for optimal results. Fig.5 and 6 illustrate the results on four datasets for univariate and multivariate forecasting, respectively. It is observed that as $\lambda$ increases from 00 to $0.5$, the prediction accuracy increases significantly and peaks at $\lambda=0.5$. This suggests that the ratio of self-supervised and supervised tasks is optimal at this point. However, when $\lambda>0.5$, the performance tends to decrease. The reason is that the proportion of self-learning is too low, which influences the performance of the model. These results verify the advantage of jointly optimizing the supervised and self-supervised tasks.
@@ 5.4.2 The Impact of $m$
In this section, we evaluate the influence of momentum coefficient on the forecasting performance. Fig. 7 and 8 illustrates the effect of different values of $m\in\{0,0.9,0.99,0.999\}$ on the varying sizes of the look-back window for ETTh1 and ETTh2 datasets. We observe that as $m$ gets closer to $1.0$, the experimental results are consistently improved, and the best results are achieved when $m=0.999$, which is consistent with the case of CV[11]. The trends in MAE and MSE values are attributed to the rapid and abrupt changes in parameters due to the 0.9 momentum update ratio. As a result, the predictive performance of the model deteriorates (worse than 0) because the rapidly changing parameters cannot effectively adapt to the transition in identifying reversal points. Therefore, momentum contrasting with slowly updating encoder parameter in our model allows for a more stable learning process, which is especially essential in time series forecasting tasks where maintaining consistency and continuity in the learned representations is crucial.
Refer to caption (a) MSE on ETTh1
Refer to caption (b) MAE on ETTh1
Refer to caption (c) MSE on ETTh2
Refer to caption (d) MAE on ETTh2
@@
Refer to caption (a) MSE on ETTh1
Refer to caption (b) MAE on ETTh1
Refer to caption (c) MSE on ETTh2
Refer to caption (d) MAE on ETTh2
@@ 5.5 Visualization Analysis
To further evaluate the forecasting quality of the proposed DE-TSMCL, we present an additional visualization analysis. We select time serie in ETTh2 to visualize the differences between ground truth and values predicted by DE-TSMCL as shown in Fig.9. Prediction results of TS2Vec[37], which performs best among baselines, are also plotted as a comparison. To better demonstrate the robustness of our model to time series distribution shifts and anomalous changes, we have zoomed in Fig. 9(c) and Fig. 9(d). In other words, Fig. 9(a) and Fig. 9(b) are sections of Fig. 9(c) and Fig. 9(d).
Fig. 10(a) and Fig. 10(b) are the scatter plots of 9(c) and Fig. 9(d). The following conclusions are obvious from the figures. We observe that (1) The ground truth curve is very irregular and it fluctuates wildly at some points, however, DE-TSMCL can fit the curve, and provide stable and accurate forecasting performance. (2) Compared with the SOTA model(i.e., TS2Vec), DE-TSMCL presents a more precise modeling ability. This indicates the effectiveness and efficiency of our proposed model by leveraging the integration of self-supervised task and supervised task, momentum contrastive learning for time series representation and prediction.
Moreover, we find that the DE-TSMCL model exhibits considerable robustness towards abnormal variations in time series. It is worthwhile to highlight that in areas of unusual transition, e.g., $t\in[300,450]$ on the ETTh2 dataset show on Fig. 9(a) and Fig. 9(b), the same conclusion can be drawn by observing Fig. 10(a). DE-TSMCL demonstrates a remarkable enhancement in forecasting performance. In addition, DE-TSMCL exhibits good performance even in continuously changing areas for a long interval, e.g., $t\in[400,500]$ at a look-back window length of 720 on the ETTh2 dataset show on Fig. 9(c) and Fig. 9(d). The same conclusion can be drawn by observing Fig. 10(b). This is attributed to the efficient momentum update module and additional supervised task. The results also verify that our proposed model is sensitive to dramatically varying peaks or troughs.
Refer to caption (a) Visualizing the DE-TSMCL part prediction.
Refer to caption (b) Visualizing the TS2vec part prediction.
Refer to caption (c) Visualizing the DE-TSMCL prediction (l=720).
Refer to caption (d) Visualizing the TS2vec prediction (l=720).
@@
Refer to caption (a) Complete scatter plot.
Refer to caption (b) Partial scatter plot.
@@
5.6 Data Augmentation Analysis(RQ4)
In this section, we provide a more comprehensive understanding of the effect of data augmentation on the model. In the design of our DE-TSMCL, data augmentation is performed after the projection layer for the teacher network, while before the projection layer for the student network. This hybrid approach differs from traditional distillation methods, which commonly adopt the same architecture for both the teacher and student networks. TABLES 7 and 8 report the comparison results for univariate and multivariate time serial forecasting on four datasets. In particular, DE-TSMCL-pre utilizes augmentation before the projection layer for both the teacher and student networks, while DE-TSMCL-after applies augmentation after the projection layer for both networks. We observe that the hybrid design (DE-TSMCL) outperforms the other design in most cases across the evaluated datasets. The main reason is that applying data augmentation after the projection layer for the teacher network helps to introduce diversity and variability into the augmented samples. This can be beneficial for the teacher network’s training, as it encourages robustness and better generalization by exposing it to a more diverse range of augmented samples. On the other hand, implementing data augmentation before the projection layer for the student network allows the model to learn from the augmented samples with increased variability and complexity. This can help the student network to better capture and understand the augmented data, potentially leading to improved performance in terms of accuracy and generalization. Therefore, this evaluation provides insights into the optimal utilization of augmentation techniques in the distillation process.
Table 7: The effect of data augmentation for univariate time series forecasting. Bold represents the best performance, and Underline highlights the second best results.
@@ 5.6 Data Augmentation Analysis(RQ4)
Table 8: The effect of of data augmentation for multivariate tiem series forecasting. Bold represents the best performance, and Underline highlights the second best results.
@@ 6 Conclusion
In this paper, we propose a Distillation Enhanced Time Series Network with Momentum Contrastive Learning(DE-TSMCL). Under the DE-TSMCL framework, we introduce the knowledge distillation between models that relay on overlapping subseries to represent the time series data for the forecasting task. Furthermore, we combine the benefits of two types of tasks to enrich the expressiveness of the prediction model. We conduct extensive experiments on four real-world datasets, and the experimental results demonstrate that DE-TSMCL is more effective compared with other SOTA models across diverse scenarios. Moving forward, we will concentrate on exploring additional convolutional techniques, including sparse convolution in 3D point clouds and novel CNN architectures, and try a more novel loss function approach such as Ranking-Based Cross-Entropy Loss and Triplet Loss, we plan to explore our framework for other time-series analysis tasks.

### v2

@@ 5.4 Hyperparameter Analysis(RQ3)
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) Electricity
@@
(a) ETTh1
(b) ETTh2
(c) ETTm1
(d) Electricity
@@ 5.4.1 The Impact of $\lambda$
As our model jointly optimizes the supervised and self-supervised tasks with hyperparameter $\lambda$ in Eq.(14), we first explore the effect of $\lambda$ on the model performance. The value of $\lambda$ determines the importance of the self-supervised task relative to the supervised task, we tune it in $\{0,0.1,0.25,0.33,0.5,0.66,0.8,0.9,1.0\}$ to find the best balance for optimal results. Fig.5 and 6 illustrate the results on four datasets for univariate and multivariate forecasting, respectively. It is observed that as $\lambda$ increases from 00 to $0.5$, the prediction accuracy increases significantly and peaks at $\lambda=0.5$. This suggests that the ratio of self-supervised and supervised tasks is optimal at this point. However, when $\lambda>0.5$, the performance tends to decrease. The low proportion of self-supervised learning has an impact on the model’s performance, which is the reason for the observed results. These findings confirm the advantage of jointly optimizing both supervised and self-supervised tasks.
@@ 5.4.2 The Impact of $m$
In this section, we evaluate the influence of momentum coefficient on the forecasting performance. Figure 7 and 8 depict the impact of different values of $m\in{0,0.9,0.99,0.999}$ on the varying sizes of prediction length for the ETTh1 and ETTh2 datasets. It can be observed that as $m$ approaches 1.0, the experimental results consistently improve. The best results are achieved when $m=0.999$, which aligns with the findings in the case of CV[12]. The reason is that our model incorporates momentum contrast with a slowly updating encoder parameter, which promotes a more stable learning process. This stability is especially crucial in time series forecasting tasks, where maintaining consistency and continuity in the learned representations is essential for accurate predictions.
(a) MSE on ETTh1
(b) MAE on ETTh1
(c) MSE on ETTh2
(d) MAE on ETTh2
@@
(a) MSE on ETTh1
(b) MAE on ETTh1
(c) MSE on ETTh2
(d) MAE on ETTh2
@@ 5.5 Visualization Analysis
To further evaluate the forecasting quality of the proposed DE-TSMCL, we present an additional visualization analysis. We select time serie in ETTh2 to visualize the differences between ground truth and values predicted by DE-TSMCL as shown in Fig.9. Additionally, we plot the prediction results of TS2Vec, which performs the best among the baseline methods, for comparison. In order to demonstrate the robustness of our model to shifts in time series distribution and anomalous changes, we have zoomed in on specific sections of the visualizations. Figure 9(c) and Figure 9(d) provide a closer look at these selected sections. To further analyze and compare the differences, we create scatter plots of Figure 9(a) and Figure 9(c), which are shown in Figure 10.
Our observations from the visual analysis are as follows:(1) The ground truth curve exhibits irregular and wild fluctuations at certain points. However, DE-TSMCL demonstrates the ability to fit the curve accurately and provides stable and precise forecasting performance. This indicates that our model can effectively capture the complex dynamics and fluctuations present in the time series data. (2) When comparing DE-TSMCL with the state-of-the-art model, TS2Vec, we observe that DE-TSMCL showcases a more precise modeling ability. This suggests that our proposed model, which integrates self-supervised and supervised tasks with momentum contrastive learning, is effective and efficient in terms of time series representation and prediction.
(a) Visualizing the DE-TSMCL prediction (P=720).
(b) Visualizing the TS2vec prediction (P=720).
(c) Visualizing the DE-TSMCL part prediction.
(d) Visualizing the TS2vec part prediction.
@@
(a) Complete scatter plot.
(b) Partial scatter plot.
@@
Table 8: The effect of data augmentation for univariate time series forecasting.
@@
Table 9: The effect of of data augmentation for multivariate time series forecasting.
@@
5.6 Data Augmentation Analysis(RQ4)
In this section, we provide a more comprehensive understanding of the effect of data augmentation on the model. In the design of our DE-TSMCL, data augmentation is performed after the projection layer for the teacher network, while before the projection layer for the student network. This hybrid approach differs from traditional distillation methods, which commonly adopt the same architecture for both the teacher and student networks. Tables 8 and 9 report the comparison results for univariate and multivariate time serial forecasting on four datasets. Bold represents the best performance, underline represents the second best performance. In particular, DE-TSMCL-pre utilizes augmentation before the projection layer for both the teacher and student networks, DE-TSMCL-after applies augmentation after the projection layer for both networks, and DE-TSMCL-reverse applies data augmentation before the projection layer for the teacher network, while after the projection layer for the student network.
We observe that DE-TSMCL outperforms the other design in most cases across the evaluated datasets. The main reason is that applying data augmentation after the projection layer for the teacher network helps to introduce diversity and variability into the augmented samples. This can be beneficial for the teacher network’s training, as it encourages robustness and better generalization by exposing it to a more diverse range of augmented samples. On the other hand, implementing data augmentation before the projection layer for the student network allows the model to learn from the augmented samples with increased variability and complexity. This can help the student network to better capture and understand the augmented data, potentially leading to improved performance in terms of accuracy and generalization. Therefore, this evaluation provides insights into the optimal utilization of augmentation techniques in the distillation process.
5.7 Statistical Significance Assessing
In this section, we have incorporated the use of the Kolmogorov-Smirnov (K-S) test to assess the distributional similarity between the input and output sequences of various models. This statistical method helps determine whether the prediction results align with the distribution of the prediction sequence. The experiment is specifically conducted on the ETTm1 dataset with a fixed step size of 96, and the results of p-value are presented in Table10. The p-value in the context of K-S test represents the probability of observing a test statistic as extreme as the one calculated from the data. In the time series forecasting, a p-value less than a predetermined significance level (e.g 0.01) suggests that the observed difference between the input and output sequences is statistically significant. In our study, the p-values obtained from the K-S test are compared to the significance level of 0.01. When the p-values for the Transformer, Informer, and TS2Vec models are below 0.01, it suggests that the input and output sequences likely originate from different distributions. On the other hand, our proposed model yields a higher p-value, it indicates that the input and output sequences are likely to be from the same distribution. This finding further supports the superiority of our model in accurately capturing the distributional characteristics of the prediction sequence.
Table 10: The K-S test results of different methods for univariate forecasting on ETTm1 dataset.
dataset P Transformer Informer TS2Vec DE-TSMCL
24 0.011 0.01 0.014 0.035
48 0.0097 0.0078 0.0074 0.029
ETTm1 96 0.0061 0.0039 0.0051 0.025
288 0.0022 0.0019 0.0037 0.016
672 0.0023 0.0016 0.0028 0.0094
@@ 6 Conclusion
In this paper, we propose a novel framework called Distillation Enhanced Time Series Network with Momentum Contrastive Learning (DE-TSMCL). DE-TSMCL incorporates knowledge distillation between models that utilize overlapping sub-series to represent time series data in forecasting tasks. Additionally, we leverage the advantages of two types of tasks: adaptive supervised task and self-supervised task, within the DE-TSMCL framework. This combination allows us to enhance the model’s representation and improve its forecasting performance. We conduct extensive experiments on five real-world datasets to evaluate the performance of DE-TSMCL. The results demonstrate that DE-TSMCL outperforms other state-of-the-art models in various scenarios. Notably, on the ETTm1 dataset, DE-TSMCL achieves a significant 24.2% improvement in Mean Squared Error (MSE) and a 14.7% improvement in Mean Absolute Error (MAE). Similarly, on the ETTh1 dataset, DE-TSMCL achieves substantial gains, with a 14.2% improvement in MSE and a 10.6% improvement in MAE.
We acknowledge that our proposed method has certain limitations. These limitations provide opportunities for future improvement and research. Some of the limitations of our proposed method include: (1) Scalability: Our method’s performance may be affected when dealing with extremely large-scale time series datasets. The computational and memory requirements of the model may become a limiting factor. To address this, we plan to explore techniques such as model parallelism, distributed computing, or more efficient architectures to improve scalability. (2) Generalization to diverse domains: Although our method demonstrates promising results in the specific domain we evaluated, its generalizability to diverse domains and datasets needs further investigation. We aim to conduct experiments on a wider range of datasets from different domains to assess the robustness and adaptability of our method.
In the future, our focus will be on exploring additional convolutional techniques, such as sparse convolution in 3D point clouds, as well as novel CNN architectures to enhance the capabilities of our model. We also aim to investigate more novel loss function approaches, such as Ranking-Based Cross-Entropy Loss and Triplet Loss to enhance the training process and improve the performance of the model. Furthermore, we plan to extend the applicability of our framework to other time-series analysis tasks.